# Don’t drop your samples! Coherence-aware training benefits Conditional diffusion

### v1

@@ Abstract
Baseline
Filtered
Weighted
CAD (Ours)
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Prompt: A vibrant, multicolored furry wolf with neon highlights playing an electric guitar on stage; trending on artstation
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Prompt: A shanty version of Tokyo, new rustic style, bold colors with all colors palette, video game, genshin, tribe, fantasy, overwatch
Figure 1: Images generated with a RIN model trained with different handling of the misalignment between the image and its associated text at training. Compared to doing nothing (baseline), removing misaligned samples (filtering) or weighting the loss accordingly (weighted), our Coherence-Aware Diffusion training (CAD) generates more visually pleasing images while better adhering to the prompt’s subject.
@@ 1 Introduction
Instead of discarding the vast majority of training samples, in this work, we leverage them to learn simultaneously conditional and unconditional distributions. Specifically, we introduce a novel approach that estimates what we call the coherence score, which evaluates how well the condition corresponds to its associated image. We incorporate this coherence score into the training process by embedding it into a latent vector, which is subsequently merged with the condition. This additional information enables the diffusion model to determine the extent to which the condition should influence the generation of a target image. During inference, our method has the flexibility to take as input the coherence score, thereby allowing users to vary the impact of the condition on the generation process, as illustrated in Figure 1. In addition, to further improve the generated image quality, we refine the Classifier-Free-Guidance method (CFG) introduced in [22] by leveraging the gap between high and low coherence scores.
@@ 3.1 Conditional Diffusion Models
Method $\omega$ FID $\downarrow$ CLIPScore $\uparrow$ P $\uparrow$ R $\uparrow$ D $\uparrow$ C $\uparrow$
@@ 4.1 Experimental setup and Metrics
For text-conditional image generation, we use a modified version of RIN [27]. To map the text to an embedding space, we use a frozen FLAN-T5 XL [9]. We then map the embedding with 2 self-attention transformer layers initialized with LayerScale [56]. We finally add the conditioning to the latent branch of RIN at each RIN Block with a cross-attention layer. The model has 188M parameters and we train it for 240K steps. We train these models on a mix of datasets composed of CC12M [5] and LAION Aesthetics 6+ [49]. To estimate the coherence score, we use MetaCLIP H/14 [58] that we then bin into 8 equally distributed discrete bins. We then use the normalized index between 0 and 1 as the coherence score. We compare our method to 3 baselines: "Baseline" is a model where we just train without coherence, "Filtered" corresponds to a model where we discard the 3 less coherent bins, and "Weighted" corresponds to a model where we weight the loss of the model by the normalized coherence of the sample. When using Coherent Aware prompting, we sample a random set of characters that we use with coherence-score of zero as the negative prompt.
@@
4.2 Analysis
Coherence conditioning.
Qualitatively, for text-to-image generation, we prompt the model with varying coherence scores from 0 to 1 and display results in Figure 2(a). We observe that when the coherence increases, the outputs are close to the prompt. For instance, in the bottom figure, the generated image displays an avocado armchair, where avocado and armchair are successfully mixed. Even a more complex prompt, like the raccoon at the top, follows closely the textual description. The raccoon does wear an astronaut suit and is looking through the window at a starry night. Similarly, as the coherence decreases, the images start to diverge from the original prompt. The avocado chair starts to first lose the "avocado" traits until there is only a chair and at the end an object that does not look like an avocado or a chair. At the top, we first lose the window, then the raccoon. We note that contrary to class conditional (as seen below), we do not converge to a totally random image. Instead, some features from the prompt are preserved, such as the astronaut suit and the starry night. This is highly linked to the CLIP network biases, which may pay less attention to less salient parts of an image such as the background, and are more sensitive to the main subject.
@@ 4.2 Analysis
Figure 5: Impact of coherence on the model. Top: (a) Impact on FID and Accuracy of prompting a model (CAD on ImageNet with $\beta=0.5$). (b) FID vs Accuracy (using CA-CFG). We vary the guidance rate from 0 to 25. (c) Impact of prompting with different coherence scores on image generation. Low coherence indicates convergence towards an unconditional model. Bottom: (d) Quantitative results for class-conditional image generation. Our coherence aware diffusion (CAD) is compared to a baseline model and a training set filtering strategy for different levels of label noise $\beta$. We show that CAD achieves higher fidelity and better accuracy. (e) TSNE of a mixed embedding of the class label and the coherence score on CIFAR-10. Each color denotes a class and the transparency shows the coherence level: the more transparent, the less confidence.
@@ 5 Conclusions
We proposed a novel method for training conditional diffusion models with additional coherence information. By incorporating coherence scores into the conditioning process, our approach allows the model to dynamically adjust its reliance on the conditioning. We also extend the classifier-free guidance, enabling the derivation of conditional and unconditional models without the need for dropout during training. We have demonstrated that our method, called condition-aware diffusion (CAD), produces more diverse and realistic samples on various conditional generation tasks, including classification on CIFAR10, ImageNet and semantic segmentation on ADE20k.
@@ Appendix A CAD architecture
A.1 RIN architecture for class conditional CAD
@@
A.2 CAD with text conditioning
To enable text-conditioned image generation, we propose a modification to the RIN architecture, coined Text RIN Block (see Figure 8). First, the text tokens are mapped with the coherence with 2x self-attention layers initialized with LayerScale [55] (top part in the figure) and 16 registers [10]. This mapping is the same for every Text RIN block. Note that unlike the class conditional RIN block (Figure 7), in our proposed Text RIN Block, the latent branch contains only the latents; there are no concatenated tokens anymore. Instead, the mapped text tokens, coherence, and timestep embeddings provide information to the latent branch with a cross-attention layer at the beginning of the Text RIN block. Then, the rest of the architecture is the same as the RIN block.
Similarly, during inference our architecture takes as input a noisy version of the image, the timestep and coherence tokens, and a text conditioning that is mapped with a frozen FLAN T5 XL encoder. It then predicts the noise that has been added to the clean image. Output latents from one denoising step are also forwarded to the next denoising step.
Figure 8: Architecture of the proposed Text RIN Block used in CAD.
@@ Appendix F Additional Qualitative Results
Figure 12: Samples from our CAD model at 512 resolution with associated caption
@@
Figure 13: Samples from our CAD model at 512 resolution
Figure 14: Samples from our CAD model at 512 resolution
Figure 15: Samples from our CAD model at 512 resolution
Figure 16: Samples from our CAD model at 512 resolution

### v2

@@ Abstract
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 1: Images generated from our model, CAD. Our model showcase high visual quality, aesthetics and prompt following.
@@ 1 Introduction
Instead of discarding the vast majority of training samples, in this work, we leverage them to learn simultaneously conditional and unconditional distributions. Specifically, we introduce a novel approach that estimates what we call the coherence score, which evaluates how well the condition corresponds to its associated image. We incorporate this coherence score into the training process by embedding it into a latent vector, which is subsequently merged with the condition. This additional information enables the diffusion model to determine the extent to which the condition should influence the generation of a target image. During inference, our method has the flexibility to take as input the coherence score, thereby allowing users to vary the impact of the condition on the generation process, as illustrated in Figure LABEL:fig:teaser. In addition, to further improve the generated image quality, we refine the Classifier-Free-Guidance method (CFG) introduced in [21] by leveraging the gap between high and low coherence scores.
@@ 3 Coherence-Aware Diffusion (CAD)
Baseline
Filtered
Weighted
CAD (Ours)
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
A vibrant, multicolored furry wolf with neon highlights playing an electric guitar on stage; trending on artstation
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
A shanty version of Tokyo, new rustic style, bold colors with all colors palette, video game, genshin, tribe, fantasy, overwatch
Figure 3: Images generated with a TextRIN trained with different handling of the misalignment between the image and its associated text at training. Compared to doing nothing (baseline), removing misaligned samples (filtering) or weighting the loss (weighted), our Coherence-Aware Diffusion training (CAD) generates more visually pleasing images while better adhering to the prompt.
@@ 3.1 Conditional Diffusion Models
Figure 4: Text RIN Block. Architecture of the proposed Text RIN Block used in CAD. We include a cross attention from the text to the latent branch of the RIN block.
@@
Method $\omega$ $\text{FID}_{\text{CLIP}}$downarrow CLIPScore $\uparrow$ $\text{P}_{\text{CLIP}}$uparrow $\text{R}_{\text{CLIP}}$uparrow $\text{D}_{\text{CLIP}}$uparrow $\text{C}_{\text{CLIP}}$uparrow
@@ 4.1 Experimental setup and Metrics
For text-conditional image generation, we use a modified version of RIN [26] (See Figure 4). To map the text to an embedding space, we use a frozen FLAN-T5 XL [9]. We then map the embedding with 2 self-attention transformer layers initialized with LayerScale [55] and 16 registers. The text tokens are mapped with the coherence score using these layers, which is the same for every Text RIN block. Unlike the class conditional RIN block, the latent branch contains only the latents without concatenated tokens. Instead, the mapped text tokens, coherence, and timestep embeddings provide information to the latent branch with a cross-attention layer at the beginning of each Text RIN block. The rest of the architecture follows the standard RIN block design. The CAD-S model has 188M parameters and we train it for 240K steps. We train these models on a mix of datasets composed of CC12M [5] and LAION Aesthetics 6+ [48]. We also train a 332M parameters version for 500k steps that we name CAD-B. We also finetune it for 200K steps at 512px resolution. To estimate the coherence score, we use MetaCLIP H/14 [57] that we then bin into 8 equally distributed discrete bins. We then use the normalized index between 0 and 1 as the coherence score. We compare our method to 3 baselines: "Baseline" is a model where we just train without coherence, "Filtered" corresponds to a model where we discard the 3 less coherent bins, and "Weighted" corresponds to a model where we weight the loss of the model by the normalized coherence of the sample. When using Coherent Aware prompting, we sample a random set of characters that we use with coherence-score of zero as the negative prompt.
@@
Figure 7: Impact of coherence on the model. Top: (a) Impact on FID and Accuracy of prompting a model (CAD on ImageNet with $\beta=0.5$). (b) FID vs Accuracy (using CA-CFG). We vary the guidance rate from 0 to 25. (c) Impact of prompting with different coherence scores on image generation. Low coherence indicates convergence towards an unconditional model. Bottom: (d) Quantitative results for class-conditional image generation. Our coherence aware diffusion (CAD) is compared to a baseline model and a training set filtering strategy for different levels of label noise $\beta$. We show that CAD achieves higher fidelity and better accuracy. (e) TSNE of a mixed embedding of the class label and the coherence score on CIFAR-10. Each color denotes a class and the transparency shows the coherence level: the more transparent, the less coherence.
4.2 Analysis
Coherence conditioning.
Qualitatively, for text-to-image generation, we prompt the model with varying coherence scores from 0 to 1 and display results in Figure 2(a). We observe that when the coherence increases, the outputs are close to the prompt. For instance, in the bottom figure, the generated image displays an avocado armchair, where avocado and armchair are successfully mixed. Even a more complex prompt, like the raccoon at the top, follows closely the textual description. The raccoon does wear an astronaut suit and is looking through the window at a starry night. Similarly, as the coherence decreases, the images start to diverge from the original prompt. The avocado chair starts to first lose the "avocado" traits until there is only a chair and at the end an object that does not look like an avocado or a chair. At the top, we first lose the window, then the raccoon. We note that contrary to class conditional (as seen below), we do not converge to a totally random image. Instead, some features from the prompt are preserved, such as the racoon but the global structure of the image gets lost. This is highly linked to the CLIP network biases, which may pay less attention to less salient parts of an image such as the background, and are more sensitive to the main subject.
@@ 5 Conclusions
We proposed a novel method for training conditional diffusion models with additional coherence information. By incorporating coherence scores into the conditioning process, our approach allows the model to dynamically adjust its reliance on the conditioning. We also extend the classifier-free guidance, enabling the derivation of conditional and unconditional models without the need for dropout during training. We have demonstrated that our method, called coherence-aware diffusion (CAD), produces more diverse and realistic samples on various conditional generation tasks, including classification on CIFAR10, ImageNet and semantic segmentation on ADE20k.
@@ Appendix A CAD architecture
A.1 RIN architecture for class conditional with coherence
@@ Appendix F Additional Qualitative Results
Figure 13: Samples from our CAD-B model at 512 resolution with associated caption
@@
Figure 14: Samples from our CAD-B model at 512 resolution
Figure 15: Samples from our CAD-B model at 512 resolution
Figure 16: Samples from our CAD-B model at 512 resolution
Figure 17: Samples from our CAD-B model at 512 resolution